OpenStudy (ebonyserrano):

The sum of the number divided by 4 and 8 is equal to 6. c for the unknown number

OpenStudy (ebonyserrano):

@Will.H

OpenStudy (will.h):

The sum means addition a number means c divided by 4 means 4/c equal to 6 = 6 4/c + 8 = 6

OpenStudy (ebonyserrano):

is the c suppose to be first or no? because it says it is incorrect

OpenStudy (will.h):

right c/4 + 8 = 6

OpenStudy (ebonyserrano):

thank you so much!!!!!!!

OpenStudy (will.h):

my pleasure :)

OpenStudy (will.h):

you sure you done this time? haha

OpenStudy (ebonyserrano):

I have more but I'll just ask someone else cause i don't want to bother lol

OpenStudy (will.h):

that's sweet haha but no i will help you i will let you know when i have to go

OpenStudy (ebonyserrano):

YOU ARE MY LIFESAVER LOL

OpenStudy (will.h):

Lol